You are assigned to manage a 12-month project estimated to cost $36,000 at the end of the project. Today is the end of month 3 and you are 20% complete. According to the finance Department you have spent $7500 to date.

Calculate the PV, EV, AC and the CPI.
Answer #1

Budget (Planned value) = 36000 for 12 months

  • Planned Value (PV) 3 months = 36000*3/12 = $9000
  • Earned value (EV) 3 months = 20%*36000 = $7200
  • Actual cost (AC) = $7500
  • CPI = EV / AC = 7200 / 7500 = 24 / 25 = 0.96